๐šƒ๐š‘๐šŽ ๐š–๐šŠ๐š๐š—๐šŽ๐š๐š’๐šŒ ๐š๐š•๐šž๐šก ๐š˜๐š 500๐š ๐š‹ ๐š™๐šŠ๐šœ๐šœ๐š’๐š—๐š ๐š๐š‘๐š›๐š˜๐šž๐š๐š‘ ๐šŠ200๐š๐šž๐š›๐š— ๐šŒ๐š˜๐š’๐š•๐š’๐šœ ๐šœ๐™ด๐š ๐š’๐š— 20ร—10๐š™๐š˜๐š ๐šŽ๐š›-3 ๐šœ๐šŽ๐šŒ๐š˜๐š—๐š๐šœย  ๐šƒ๐š‘๐šŽ ๐šŠ๐šŸ๐šŽ๐š›๐šŠ๐š๐šŽ ๐šŽ. ๐™ผ. ๐™ต ๐š’๐š—๐š๐šž๐šŒ๐šŽ๐š ๐š’๐š— ๐š๐š‘๐šŽ ๐šŒ๐š˜๐š’๐š• ๐š’๐š— ๐šŸ๐š˜๐š•๐š ๐š’๐šœ

Grade:12

2 Answers

Vikas TU
3 years ago
Dear studentย 
e = -N (d phi / dt)ย 
d phi = phi 2 - phi 1ย 
dt = 2*10^-3ย 
N = 200ย 
calculating e = 10 volt .ย 
Which is magnitude of e.ย 
Arun
3 years ago
Dear student
 
Emf=N?t??
= 20×10 
−3
2×500×10 
−6 ×200
?=10V/m 
2
Hope you understand
Thanks and regards
Arun